OpenStudy (anonymous):

how do i solve 12x^2=36. Give both exact solutions and approximate solution to the nearest hundredth.

sam (.sam.):

You don't need quadratic formula here, 12x^2=36 x^2=3 \[x= \pm \sqrt{3}\]

OpenStudy (anonymous):

so basically how do i do this? do i just simplify?

sam (.sam.):

\[ 12x^2=36\] divide both sides by 12 \[ x^2=3\] square root both sides x=±√3

OpenStudy (anonymous):

so what if its (x+3)^2=81

sam (.sam.):

then square root both sides first (x+3)^2=81 \[x+3=\pm \sqrt{81}\]

OpenStudy (anonymous):

oh okay so its ±√81?

sam (.sam.):

\[x+3=\pm \sqrt{81}\] \[x+3=\pm 9\] ------------------------------------ x+3=9 and x+3=-9 x=6 x=-12

OpenStudy (jiteshmeghwal9):

If there is square of x at L.H.S then after taking it to other side it will be sqrt of 36 so ,\[12x=\sqrt{36}\]\[12x=6\]\[x=\frac{6}{12} \implies \frac{1}{2}\]

OpenStudy (jiteshmeghwal9):